Determine pH: 33.9 mL of 5.88 x 10-3 M HClO4 is mixed with 207 mL of 2.17 x 10-2 M SrO


1
Expert's answer
2021-12-08T01:23:01-0500

pOH of the base = 14 - 13.70 = .3

[OH] = 10^-.3 = .5012M

Since both of these are strong acids, we just need to find which will be left over after the reaction

mole of acid: .335×1.358 = .455mol

mole of base: .585×.5012  = .2932mol base

.455 - .2932 = .1618 mole of acid left over

New [H+] = .1618/(.335+.585) = .176M

pH = -log(.176) =0.755